a local hair salon receives a shipment of 84 bottles of hair conditioner to use and sell to customers. The two types of conditioners received are type A, which is used for straight hair, and type B which is used for curly hair. Type A costs $6.50 per bottle and type B cocsts $8.25 per bottle. The salon paid a total of $588 for the conditioner. How many of each type of conditioner did the local salon buy?
Answer by mananth(16946)

a local hair salon receives a shipment of 84 bottles of hair conditioner to use and sell to customers. The two types of conditioners received are type A, which is used for straight hair, and type B which is used for curly hair. Type A costs $6.50 per bottle and type B cocsts $8.25 per bottle. The salon paid a total of $588 for the conditioner. How many of each type of conditioner did the local salon buy?
Let type a be x
Type a------------------------- Type b-------------- total
x--------------------------------84-x----------------84
$6.5-----------------------------$8.25---------------$588

6.5x+(84-x)*8.25= 588
6.5x +693-8.25x = 588
-1.75x=-693-588
x= -105/1.75
x= 60
Type a is 60 numbers
Type b will be 84-60= 24 type b